OpenStudy (anonymous):

As you can see the biker and his friend are having a hard time getting up the hill. They would like to get to the top safely. Two ordered pairs on the hill are given as (6, 3) and (8, 6). Part 1: Slope (m) = change in y over change in x. y sub two minus y sub one all over x sub two minus x sub one Part 2: Now find the equation using the point slope formula: Point Slope Formula = y – y1 = m (x – x1) Part 3: Give an example of an equation for a line parallel to your answer from Part 2 Part 4: Give an example of an equation for a line perpendicular to your answer from Part

OpenStudy (anonymous):

if these are the points (6, 3) and (8, 6). and changed in y y2-y1 slope m= -------------- =-------------- changed in x x2=x1 can you start solving them?

OpenStudy (anonymous):

hint (6, 3)=(x1,y1) and (8, 6)=(x2,y2)

OpenStudy (anonymous):

okay i got 6-3/8-6 = 3/2 now what?

OpenStudy (anonymous):

ok now that you have slope m=3/2, use this in part 2) Point Slope Formula = y – y1 = m (x – x1) form that one, by using one of the points on (6, 3)=(x1,y1) and (8, 6)=(x2,y2)

OpenStudy (anonymous):

are you there @ IssiahZeek